OpenStudy (anonymous):

the total area of a cube is144. what is its volume? a.24 cu units b. 96 Square root of 6 cu units c.48 square root of 6 cu units d. 144 cu units

OpenStudy (anonymous):

alright so a cube has 6 sides, and the area of one square is x*x. (you have six of these) So, 6x^2=144 x=4.898 x^3=volume so (4.898)^3=48 root 6

OpenStudy (raden):

6x^2 = 144 x^2 = 144/6 = 24 x = sqrt(24) = 2sqrt(6) volume = side * side * side = 2sqrt(6) * 2sqrt(6) * 2sqrt(6) = 8 * 6 * sqrt(6) = 48sqrt(6)